A 1.0 H.P. motor pumps out water from a well of depth 20m and fills a water tank of volume 2238 liters at a height of 10 m from the ground. The running time of the motor to fill the empty water tank is (g = 10ms)?

Grade:12th pass

3 Answers

Vikas TU
7 years ago
Mass of water from well => density x volume => 1*2238 gms or 2.238 Kg.
Work  = mgh => 2.238*10*20 => 447.6 J
 
For height 10 m from the ground= >  2.238*10*10 => 223.8 J
Net Energy required = > 447.6 – 223.8 => 223.8 J
1 HP =746 w = 746 J
time required => 223.8/746 => 0.3 seconds.
jitender
6 years ago
P=1hp = 746wattd=20m ,h=10m ~> H=10+20= 30mVol=2238ltr{1ltr=1Kg : 1ltr=10^-3m^3...;density=10^3Kg/m^3And mass= vol × density}t=?P=W/tt=W/P = mgh/P = 2238×10×30/746 Dect= 900sect=900/60mint=15 min
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Mass of water from well => density x volume => 1*2238 gms or 2.238 Kg.
Work  = mgh => 2.238*10*20 => 447.6 J
For height 10 m from the ground= >  2.238*10*10 => 223.8 J
Net Energy required = > 447.6 – 223.8 => 223.8 J
1 HP =746 w = 746 J
time required => 223.8/746 => 0.3 seconds
 
Thanks and regards,
Kushagra